1/2

Christian Morgenstern, Christian Morgenstern,

Christian Morgenstern,

Varulven  Le loup-garou  The Werewolf

tradukita de Sven Collberg tradukita de R. Platteau

tradukita de Alexander Gross

En varulv lopp från fru

m.m. en vacker afton samt be-

gav sig till en bymagisters

grav och bad: “Var snäll och

konjugera.

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître

d’école Et lui dit, “ sur moi dis tout, c’est

ton rôle ! ”

A Werewolf, troubled by his na-

me, Left wife and brood one night

and came To a hidden graveyard to enlist The aid of a long-dead philolo-

gist.

mig!” Byskolläraren stod

upp på blecknamnsskyltens

mässingknopp med svar till den, som

satt på pass med tåligt korslagd tass

på tass.

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître

d’école Et lui dit, “ sur moi dis tout, c’est

ton rôle ! ”

“Oh sage, wake up, please don’t

berate me,” He howled sadly, “Just conjugate

me.” The seer arose a bit unsteady Yawned twice, wheezed once, and then was ready.

“Jag var-ulv 1:a pers. i

sing., du, han, hon, den, det

var-ulv. Så vi voro-ulv, I voren. På de voro-ulv kom ingen-

ting.”

“ - On peut à ton sujet puisque nous

y sommes Se demander encor bien des choses

en somme. Loup-garoù ? mais avant Il faut se dire : loup-garquand ?

“Well, ‘Werewolf’ is your plural

past, While ‘Waswolf’ is singularly

cast: There’s ‘Amwolf’ too, the pre-

sent tense, And ‘Iswolf,’ ‘Arewolf’ in this

same sense.”

Varulven myste glad åt

hur pers. hade följts av sing.

och plur. “Fast”, bad han, “kan det

inte ges ens en enda pers. i tempus

presens?”

Loup-garcomment ? et puis

Bien évidemment loup-garqui ? ” Toutes ces questions semblèrent au

loup bien absconses, Car en fait ce qu’il voulait c’était les

réponses.

“I know that–I’m no mental

cripple– The future form and participle Are what I crave,” the beast re-

plied. The scholar paused–again he

tried:

…  …  …

2/2

Bekänna måste vår ma-

gister, att därvidlag fanns vissa

brister. En varulv sine imperfec-

to var contradictio in ad-

jecto.

“ - Point de réponses, dit le puits de

savoir, Etre imaginaire, nul ne peut te voir

! ” A ces mots l’animal éclata tout en

larmes. “ - Oh!, même les baisers de ma dou-

ce moitié “A ‘Will-be-wolf?’ It’s just too

long: ‘Shall-be-wolf?’ ‘Has-been-wolf?’ Utterly wrong! Such words are wounds beyond

all suture– I’m sorry, but you have no fu-

ture.”

Varulven satt med tårad

lins. Han visste att han fanns

och finns. Därtill kom också fru

m.m. som även måste existera.

Ne pourrons plus me faire me sentir

tout entier. Rester un loup-garou est un sort qui

m’alarme. ” Mais comme le loup était bien poli Il remercia et puis repartit.

The Werewolf knew better–his

sons still slept At home, and homewards now

he crept, Happy, humble, without apology For such folly of philology.

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Svedan de Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Francan de R. Platteau.

Arg-825-1655 (2013-02-13 21:56:38)

Tiu ĉi francigo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .